What is 9/20+6 3/40+ 8 7/8 as a fraction
Answer by Alan3354(69443)

What is 9/20+6 3/40+ 8 7/8 as a fraction
------------
Change to fractions, no mixed numbers.
= 9/20 + 243/40 + 71/8
Find a common denominator. 40 is obvious
= 18/40 + 243/40 + 355/40
= (18 + 243 + 355)/40
= 616/40
= 77/5